# Peeking with PEAK: Sequential, Nonparametric Composite Hypothesis Tests for Means of Multiple Data Streams

### v2

@@ 3.1 Asymptotic $e$-Power/Growth Rate
Let $P$ be a distribution with mean $\mu$, i.e., $P\in\mathcal{P}(\mu)$. We define the asymptotic growth rate of $K_{t}(m)$ as $G(c,m,P)=\lim_{t\rightarrow\infty}\mathbb{E}_{P}[\log(1+\frac{(\hat{\mu}_{t-1}-m)(X-m)}{c})]=\mathbb{E}_{P}[\log(1+\frac{(\mu-m)(X_{t}-m)}{c})]$  [^2] The interchange of limits and integrals in the last equality is valid by Leibnitz’s rule..
In Lemma 1, we first establishes that the asymptotic growth rate $G(c,m,P)$ serves as a proxy for the expected nonasymptotic rate of linear growth for $l_{t}(m)$ (or the expected multiplicative factor in each time step for $K_{t}(m)$).
Lemma 1 (Asymptotic Growth Rate and $K_{t}(m)$).
For all $m\in[0,1]$, $P\in\mathcal{P}(\mu)$, and for any $\epsilon>1$, there exists a finite $t(\epsilon)\in\mathbb{N}$ such that for $t>t(\epsilon)$, the nonasymptotic growth rate $\mathbb{E}_{P}[l_{t}(m)-l_{t-1}(m)]$ satisfies the following:
$\mathbb{E}_{P}[l_{t}(m)-l_{t-1}(m)]$ $\geq G(c\epsilon,m,P).$
@@
While setting $c$ to its minimal value maximizes the asymptotic growth, we recommend setting $c=1/4+\gamma$, where $\gamma$ is a small constant such as $\gamma=0.01$ for the sake of stability.333 Consider the scenario with $c=1/4$, where the estimated mean from previous observations $\hat{\mu}_{i-1}=1$, $m=0.5$, and the next observed data point $X_{i}=0$. Then, $K_{t}(m)=0$. For any time after $t$, $K_{t}(m)=0$, meaning that the hypotheses $m=0.5$ cannot be rejected for any time in the future.
In the next lemma, Lemma 3 (proof in Appendix A.2), we establish that the worst-case growth rate of $G(c,m,P)$ is strictly positive for hypotheses $m\neq\mu$, and increases as a function of the absolute difference between null hypothesis $m$ and true mean $\mu$. Combined with the results of Lemma 1, this implies that for any $m\neq\mu$, the nonasymptotic growth rate of $l_{t}(m)$ is necessarily positive at some time $t$.
@@
[…]
@@
Finally, our test is nonparametric. When compared with the likelihood ratios of Wald, (1945), naturally, one would expect our worst-case asymptotic growth rate to deviate significantly from the best achievable parametric growth rate. However, Lemma 4 (proof in Appendix A.5) shows that our test achieves near-optimal asymptotic growth rate in the simple-vs-simple Bernoulli setting, an instance of our worst possible growth rate. Let $\mu\in(0,1)$. Then, we have:
@@
[…]
@@
Figure 2: Fig (A): Plot of $G(c,m,P)$ for Null Hypothesis $m$, $P\equiv\text{Bern}(\mu)$. Fig (B): Plot of $f(c,m,P)$ for Null Hypothesis $m$, $P\equiv\text{Bern}(\mu)$. Parameter $c$ is set as $0.26$.
We note that as $|\mu-m|$ decreases, $f(m,\mu)$ increases rapidly, meaning that $K_{t}(m)$ preserves the best-possible expected growth rate in difficult instances (i.e., null hypothesis $m$ is close to alternative hypothesis $\mu$). For instances where our growth rate demonstrates suboptimality (i.e., $m\approx 0$ or $m\approx 1$), we observe that $G(c,m,P)$ is still large (white/orange regions of Fig. 2A). Lemma 4 and Figure 2 demonstrate that for incorrect hypotheses close to the true mean, the asymptotic growth rate of our process $K_{t}(m)$, in the worst case, is near-optimal with respect to the best-achievable growth rate.
@@ 4 Testing in the Multi-Stream Setting
In this section, we generalize the results of the single stream/arm setting to $W$ streams/arms, where $W\in\mathbb{N}$, and demonstrate how to test for a general null composite hypothesis $R$ that is convex in the hypothesis space $m\in[0,1]^{W}$. Let $N_{t}(a)=\sum_{i=1}^{t}\mathbf{1}[A_{i}=a]$ and $\hat{\mu}_{t}(a)=\sum_{i=1}^{t}\mathbf{1}[A_{i}=a]X_{i}/N_{t}(a)$. We first define our product capital process and the test for the case $R=m$ below:
@@
$T_{t}(m,\alpha)=\mathbf{1}[\max_{1\leq i\leq t}E_{t}(m)\geq 1/\alpha].$ (10)
@@
The naive way to generalize our sequential tests is to simply use union bounds: we divide $\alpha$ by $W$, and reject the hypothesis at time $t$ if each individual component $K_{t}^{a}(m_{a})$ exceeds the $W/\alpha$ threshold. In Lemma 5 (proof in Appendix A.9), we establish that the product test strictly dominates the union bound test.
@@
$T^{\prime}_{t}(m,\alpha)=\prod_{a\in[W]}\mathbf{1}[\max_{1\leq i\leq t}K_{i}^{a}(m_{a})\geq W/\alpha].$
For all $t\in\mathbb{N}$, any $m\in[0,1]^{W}$ rejected by union bounding test $T^{\prime}_{t}(m,\alpha)$ is also rejected by the test $T_{t}(m,\alpha)=\mathbf{1}[\max_{1\leq i\leq t}E_{i}(m)\geq 1/\alpha]$ for any realization of the data.
@@ 4.1 Theoretical Guarantees for Joint Capital Process
Let $\pi_{t}:H^{t-1}\rightarrow\Delta^{W}$ be a sampling scheme, where $H^{t-1}=\{A_{1},X_{1},...,A_{t-1},X_{t-1}\}$. This process $E_{t}(m)$ shares the same theoretical guarantees of the single arm setting for 1) type I error control for all $\pi_{t}$, and 2) power under mild assumptions of $\pi_{t}$.
@@
For $m=\mu$, $c\geq 1/4$, for any sampling scheme $\pi_{t}$, the test $T_{t}(m,\alpha)$ has type I error probability $1-\alpha$.
Theorem 3 (proof in Appendix A.7) provides type I error control for any possible sampling scheme $\pi_{t}$, and maintains the correctness of our test. To show that $T_{t}(m,\alpha)$ is a test of power one for any $m\neq\mu$, we require a relatively mild condition as stated in Theorem 4 (proof in Appendix A.8).
@@
Intuitively, the condition of Theorem 4 states that for an incorrect hypotheses $m\neq\mu$, we must pull at least one mispecified arm infinitely often (not necessarily the same arm across data realizations) as $t\rightarrow\infty$. Simple examples of sampling schemes that satisfy this property include $\pi_{t}(a)=1/W$, or bandit sampling schemes such as $\epsilon$-greedy or action elimination (Jamieson and Nowak,, 2014).
@@ 4.2 Testing Composite Convex Hypotheses
The key benefit of our testing approach, as compared to using the betting schemes of Waudby-Smith and Ramdas, (2023), is computational tractability for composite null tests that define convex regions in the hypothesis space, $[0,1]^{W}$. To reject a region $R$ of the hypothesis space, we require that every hypothesis $m\in R$ in the region is rejected. In terms of our test, the criteria to reject a region $R$ is as follows:
$T_{t}(R,\alpha)=\mathbf{1}[\min_{1\leq i\leq t,m\in R}E_{i}(m)\geq 1/\alpha].$ (11)
For such composite tests, it is essential that finding the minima of $E_{t}(m)$ within region $R$ is computationally feasible. We note that a straightforward extension of certain testing-as-betting schemes [^4] One such scheme is AGRAPA (Appendix B.4. in Waudby-Smith and Ramdas,, 2023), which can form non-convex confidence regions. by Waudby-Smith and Ramdas, (2023) to the multi-arm setting would involve a grid search over the region $R\subset[0,1]^{W}$ to test hypothesis $R$. Because our composite test across joint means is simply the average of $K_{t}^{a}(m_{a})$, it inherits the properties described in Propostion 1, such as strict convexity and a unique global minima $\tilde{m}\in[0,1]^{W}$. An immediate consequence is that the global minima of $E_{t}(m)$ is characterized by the following proposition.
@@
For time $t$, the vector $m^{*}=[m_{1}^{*},...,m_{W}^{*}]\in[0,1]^{W}$ that minimizes $E_{t}(m)$ is (1) unique, and (2) is characterized component-wise by the following condition:
$\sum_{i=1}^{t}\sum_{a\in[W]}\mathbf{1}[A_{t}=a]\frac{2m_{a}^{*}-\hat{\mu}_{i-1}(a)-X_{i}}{c+(\hat{\mu}_{i-1}(a)-m_{a}^{*})(X_{i}-m_{a}^{*})}=0$
To find the minima within region $R$, we simply project our global minima $m^{*}$, characterized in Proposition 2. We provide two explicit examples of this projection, using the examples of threshold identification (THR) and best-arm identification (BAI).
@@
To test hypothesis $R_{\text{Thresh.}}^{0}(a^{\prime})$ at time $t$, we simply need to check if $E_{t}(m^{0})$ exceeds the $1/\alpha$ threshold, where $m^{0}$ is given by:
@@
For hypothesis $R_{\text{THR}}^{1}(a^{\prime})$ at time $t$, we only need to check if $E_{t}(m^{1})$ exceeds the $1/\alpha$ threshold, where $m^{1}$ is defined as:
@@
The vectors $m^{0}$ and $m^{1}$ are the projections of the global minimizer $m_{a}^{*}$ for regions $R_{\text{THR}}^{0}(a^{\prime})$, $R_{\text{THR}}^{0}(a^{\prime})$ respectively, and only involve thresholding the mean of arm $a^{\prime}$ based on $\xi$. Below, we consider a slightly more complicated example, BAI, which has constraints defined between pairs of arms.
@@
For best-arm identification each arm $a$ has a single hypothesis, as defined in Equation (3). Proposition 3 provides the conditions that define the minima of $E_{t}(m)$ for region $R_{\text{BAI}}(a)$ at time $t$.
@@
[…]
@@
$\sum_{a\in W:\tilde{m}_{a}\geq q}\big{(}\gamma(m^{*},a)$
$\sum_{i=1}^{t}\mathbf{1}[A_{t}=a]\frac{2q-X_{i}-\hat{\mu}_{i-1}(a)}{c+(\hat{\mu}_{i-1}(a)-q)(X_{i}-q)}\big{)}=0,$
@@
Figure 3: Log-widths of the confidence sequences across 5000 time steps. Curves represents the average width across 30 simulations.
@@ 5 Empirical Results
Arm Type Stopping Rule $\tau_{1}$ $\tau_{2}$ $\tau_{3}$ $\tau_{4}$
Table 1: Stopping Times $\tau$ for THR. Each $\tau_{i}$ represents the first time in which we have classified $i$ arms as above or below $\xi=0.5$. We report averaged stopping times and one standard error over 100 simulated sample paths. Bolded values represent the best and second-best average run-time respectively. Across all simulations, each stopping time results in the correct conclusion (i.e. rejects the incorrect hypothesis).
In this section, we provide empirical comparisons for our test that demonstrates (i) in the single arm case, our test provides robust confidence sequences that verify the intuitions of Lemma 4, and (ii) in the multi-stream setting with adaptive sampling, our test provides improved empirical performance (i.e., smaller average stopping times, shorter runtimes) to verify the composite hypothesis tests corresponding to threshold identification (THR) and best arm identification (BAI). The results of this section suggest that our approach provides comparable performance to other testing-as-betting schemes, while improving on per-iteration runtime. For all simulations, we use $\alpha=0.05$, $c=0.26$.
5.1 Experiments in the Single Stream Setting
@@
In this setting, we compare the confidence sequences generated by our test to existing anytime-valid confidence intervals. We use the confidence sequence as a proxy for both the power and growth rate of our test. We compare our approach with four other nonparametric sequential tests that directly apply to bounded means: 1) Chernoff-based sub-Bernoulli (Sub-B) confidence sequences (Howard et al., 2021a, ),555as implemented using package confseq (Howard et al., 2021b, ) 2) empirical Bernstein (Emp. Bern.) style confidence sequences (Waudby-Smith and Ramdas, 2023, Thm. 2), 3) Hoeffding-based confidence sequences (PrPlH, Waudby-Smith and Ramdas, 2023, Prop. 1), and 4) the hedged capital process confidence sequence (Hedged, Waudby-Smith and Ramdas, 2023, Theorem 3, Eq. 26). For Hedged, we use a grid of 1000 points for all simulations. We provide further details in Appendix B.
5.2 Multiple Streams of Data, Adaptive Sampling
We test multiple streams by 1) introducing adaptive sampling schemes that vary over time, and 2) requiring $\alpha$-level guarantees across multiple parameters of interest.
Data-Generating Process. We set $W=4$. In the Bernoulli setting (high variance case), the distribution of arm $a$ is $P_{a}\equiv\text{Bern}(0.15+0.14a)$. In the Beta setting (low variance case), the distribution of arm $a$ is $P_{a}\equiv\text{Beta}(1,\frac{0.85-0.14a}{0.15+0.14a})$. In both cases, the means of each arm are identically $\mu_{a}=0.14a+0.15$.
Threshold Identification. We set threshold $\xi=0.5$. We use the sampling algorithm HDoC proposed by Kano et al.,, 2018: at time $t$, HDoC sample arm $a_{t}^{\text{HDoC}}$, defined as $a_{t}^{\text{HDoC}}=\operatornamewithlimits{arg\,max}_{a\in[K]}\hat{\mu}_{t}(a)+\sqrt{\frac{\log(t)}{2N_{t}(a)}}.$ We report the first time in which we have rejected $i$ regions as $\tau_{i}$, i.e. $\tau_{i}$ corresponds to the first time that we have labeled $i$ arms.
For both THR and BAI, Kano et al.,, 2018 and Jamieson and Nowak,, 2014 (Eq. 2) propose a stopping criteria that uses sequential tests that union bound over each arm. We denote their proposed stopping criteria as Base, and additionally test the two best performing tests/confidence sequences (Sub-B, Hedged) in the single arm case by using a similar union bound approach.
@@ 5.2 Multiple Streams of Data, Adaptive Sampling
5.3 Discussion of Results
In the single arm case, Figure 2 plots the log-width of the confidence sequences across time. The log-width of our confidence interval suggests that PEAK provides a robust testing procedure, even in the single arm case. Across any combination of variance settings and symmetry, our method performs the second-best across all tested approaches. In the high variance Bernoulli setting, Sub-B performs the best, as the sub-Bernoulli condition is tight with respect to the tested distribution. Compared to all other approaches, our test is closest in performance, corroborating the worst-case optimality results of Lemma 4. For the low-variance case, Hedged outperforms all other methods, but is closely followed by Emp. Bern. and PEAK. The fact that PEAK matches the performance of Emp. Bern., which explicitly takes the underlying variance into account, suggests that our test has benefits even in the single arm case.
In the multi-armed case, our test performs comparably to the state of the art testing-as-betting method, Hedged, and outperforms the proposed stopping critera and Sub-B. For both the THR and BAI, the stopping times for PEAK are either the lowest or second lowest on average empirically. For THR (Table 1), the performance of our approach closely mirrors its performance in the single-arm case, where PEAK performs second-best for both high and low variance settings, and Hedged and Sub-B perform best for Beta (low variance) and Bernoulli (high variance) arms respectively. For the best-arm identification problem, the testing-as-betting methods (Hedged and PEAK) drastically outperform Base and Sub-B as a stopping criteria for the best-arm identification problem. PEAK and Hedged have the smallest or second smallest average stopping times, regardless of the distribution of arms. For the high variance, Bernoulli setting, PEAK outperforms Hedged, while Hedged outperforms PEAK in the low variance Beta setting.
Runtimes for our Method. To stress the key benefits of our approach, we test the best-arm hypothesis in the Bernoulli setting across a fixed $T=2000$ time horizon (one test every two time steps), with no stopping, and report the average runtime for the fixed-horizon test over 100 simulations in Table 3. The runtimes for PEAK and Hedged are only comparable for a grid size of 100 (i.e. $10^{-2}$ fidelity) for Hedged. For larger grid sizes, such as 200 or 500 evenly spaced points ($5\times 10^{-3}$, $2\times 10^{-3}$ respectively), the runtime of Hedged is at least double that of PEAK, despite having far worse fidelity. The union bound approach of using confidence intervals with the Hedged betting scheme necessarily scales linearly with the fidelity of the grid, and becomes exponential if one were to use the approach of averaging across arms, as we do with PEAK. This demonstrates that even with a relatively small number of streams, our approach avoids fidelity issues and poor scaling of existing betting schemes, such as Hedged.

### v3

@@ 3.1 Asymptotic $e$-Power/Growth Rate
Let $P$ be a distribution with mean $\mu$, i.e., $P\in\mathcal{P}(\mu)$. We define the asymptotic growth rate of $K_{t}(m)$ as $G(c,m,P)=\lim_{t\rightarrow\infty}\mathbb{E}_{P}[\log(1+\frac{(\hat{\mu}_{t-1}-m)(X-m)}{c})]=\mathbb{E}_{P}[\log(1+\frac{(\mu-m)(X_{t}-m)}{c})]$. [^3] The interchange of limits and integrals in the last equality is valid by Leibnitz’s rule.
@@
While setting $c$ to its minimal value maximizes the asymptotic growth, we recommend setting $c=1/4+\gamma$, where $\gamma$ is a small constant such as $\gamma=0.01$ for the sake of stability.444 Consider the scenario with $c=1/4$, where the estimated mean from previous observations $\hat{\mu}_{i-1}=1$, $m=0.5$, and the next observed data point $X_{i}=0$. Then, $K_{t}(m)=0$. For any time after $t$, $K_{t}(m)=0$, meaning that the hypotheses $m=0.5$ cannot be rejected for any time in the future.
Next, in Lemma 2 (proof in Appendix A.2), we establish that the worst-case growth rate of $G(c,m,P)$ is strictly positive for hypotheses $m\neq\mu$, and increases as a function of the absolute difference between the null hypothesis $m$ and the true mean $\mu$.
@@
For any fixed $m\in[0,1],\mu\in[0,1]$, the worst-case instance $P^{*}:=\inf_{P\in\mathcal{P}(\mu)}G(c,m,P)$ is $P^{*}\equiv\text{Bern}(\mu)$.
@@
Finally, our test is nonparametric. When compared with the likelihood ratios of Wald, (1945), naturally, one would expect our worst-case asymptotic growth rate to deviate significantly from the best achievable parametric growth rate. However, Lemma 3 (proof in Appendix A.4) shows that our test achieves near-optimal asymptotic growth rate in the simple-vs-simple Bernoulli setting, an instance of our worst possible growth rate. Let $\mu\in(0,1)$. We have:
@@
When testing a Bernoulli distribution $X_{i}\sim\text{Bern}(\mu)$, the ratio of $G(c,m,P)$ to the best achievable growth rate in this simple-vs-simple sequential test is given by:
@@
Figure 1: Growth rate visualization for different ground truth $\mu$ and hypothesis $m$ combinations under Bernoulli distributions, $P\equiv\text{Bern}(\mu)$, and $c=0.26$. Left: Asymptotic growth rate of $K_{t}(m)$, $G(c,m,P)$, for null hypothesis $m$, with darker colors representing larger growth rates. Right: The ratio of $G(c,m,P)$ to the best achievable growth rate, $f(c,m,P)$, for null hypothesis $m$, with darker colors representing larger ratios.
We observe that as the distance between $\mu$ and $m$ decreases, $f(m,\mu)$ increases rapidly in Figure 1 (right), meaning that $K_{t}(m)$ preserves the best-possible expected growth rate in difficult instances (i.e., the null hypothesis $m$ is close to the alternative hypothesis $\mu$). For instances where our growth rate demonstrates suboptimality (i.e., $m\approx 0$ or $m\approx 1$), we observe that $G(c,m,P)$ is still large (white/orange regions of Fig. 1 left). Lemma 3 and Figure 1 demonstrate that for incorrect hypotheses close to the true mean, the asymptotic growth rate of our process $K_{t}(m)$, in the worst case, is near-optimal with respect to the best-achievable growth rate.
@@ 4 Testing in the Multi-Stream Setting
In this section, we generalize the results of the single-stream/arm setting to $W$ streams/arms, where $W\in\mathbb{N}$, and demonstrate how to test for a general null composite hypothesis $R$ that is convex in the hypothesis space $m\in[0,1]^{W}$. Let $N_{t}(a)=\sum_{i=1}^{t}\mathbf{1}[A_{i}=a]$ and $\hat{\mu}_{t}(a)=\sum_{i=1}^{t}\mathbf{1}[A_{i}=a]X_{i}/N_{t}(a)$. We first define our product capital process and the test for the case $R=m$ below:
@@
$T_{t}(m,\alpha)=\mathbf{1}\left[\max_{1\leq i\leq t}E_{t}(m)\geq 1/\alpha\right].$ (10)
@@
The naive way to generalize our sequential tests is to use union bounds: we divide $\alpha$ by $W$ and reject the hypothesis at time $t$ if each component $K_{t}^{a}(m_{a})$ exceeds the $W/\alpha$ threshold. In Lemma 4 (proof in Appendix A.8), we establish that the product test strictly dominates the union bound test.
@@
$T^{\prime}_{t}(m,\alpha)=\prod_{a\in[W]}\mathbf{1}\left[\max_{1\leq i\leq t}K_{i}^{a}(m_{a})\geq W/\alpha\right].$
For all $t\in\mathbb{N}$, any $m\in[0,1]^{W}$ rejected by union bounding test $T^{\prime}_{t}(m,\alpha)$ is also rejected by the test $T_{t}(m,\alpha)=\mathbf{1}\left[\max_{1\leq i\leq t}E_{i}(m)\geq 1/\alpha\right]$ for any realization of the data.
@@ 4.1 Theoretical Guarantees for Joint Capital Process
Let $\pi_{t}:H^{t-1}\rightarrow\Delta^{W}$ be a sampling scheme, where $H^{t-1}=\{A_{1},X_{1},...,A_{t-1},X_{t-1}\}$. This process $E_{t}(m)$ shares the same theoretical guarantees of the single arm setting for 1) type-I error control for all $\pi_{t}$, and 2) power under mild assumptions of $\pi_{t}$.
@@
For $m=\mu$, $c\geq 1/4$, for any sampling scheme $\pi_{t}$, the test $T_{t}(m,\alpha)$ has type-I error probability $1-\alpha$.
Theorem 3 (proof in Appendix A.6) provides type-I error control for any possible sampling scheme $\pi_{t}$, and maintains the correctness of our test. To show that $T_{t}(m,\alpha)$ is a test of power one for any $m\neq\mu$, we require a relatively mild condition as stated in Theorem 4 (proof in Appendix A.7).
@@
Intuitively, the condition of Theorem 4 states that for an incorrect hypothesis $m\neq\mu$, we must pull at least one misspecified arm infinitely often (not necessarily the same arm across data realizations) as $t\rightarrow\infty$. Simple examples of sampling schemes that satisfy this property include $\pi_{t}(a)=1/W$ or bandit sampling schemes such as $\epsilon$-greedy or action elimination (Jamieson and Nowak,, 2014).
@@ 4.2 Testing Composite Convex Hypotheses
The key benefit of our testing approach, as compared to using the betting schemes of Waudby-Smith and Ramdas, (2023), is computational tractability for composite null tests that define convex regions in the hypothesis space, $[0,1]^{W}$. To reject a region $R$ of the hypothesis space, we require that every hypothesis $m\in R$ in the region is rejected. In terms of our test, the criterion to reject a region $R$ is as follows:
$T_{t}(R,\alpha)=\mathbf{1}\left[\min_{1\leq i\leq t,m\in R}E_{i}(m)\geq 1/\alpha\right].$ (11)
For such composite tests, it is crucial to find the minima of $E_{t}(m)$ within region $R$ efficiently. We note that a straightforward extension of certain testing-by-betting schemes [^5] One such scheme is AGRAPA (Appendix B.4. in Waudby-Smith and Ramdas,, 2023), which can form non-convex confidence regions. by Waudby-Smith and Ramdas, (2023) to the multi-arm setting would involve a grid search over the region $R\subset[0,1]^{W}$ to test hypothesis $R$. Because our composite test across joint means is simply the average of $K_{t}^{a}(m_{a})$, it inherits the properties described in Proposition 1, such as strict convexity and a unique global minima $\tilde{m}\in[0,1]^{W}$. Consequently, the global minima of $E_{t}(m)$ is characterized by the following:
@@
For time $t$, the vector $m^{*}=[m_{1}^{*},...,m_{W}^{*}]\in[0,1]^{W}$ that minimizes $E_{t}(m)$ is 1) unique, and 2) characterized component-wise by the following condition:
$\sum_{i=1}^{t}\sum_{a\in[W]}\mathbf{1}[A_{t}=a]\frac{2m_{a}^{*}-\hat{\mu}_{i-1}(a)-X_{i}}{c+(\hat{\mu}_{i-1}(a)-m_{a}^{*})(X_{i}-m_{a}^{*})}=0.$
To find the minima within region $R$, we project our global minima $m^{*}$ as characterized in Proposition 2. Below, We provide two examples of this projection: threshold identification (THR) and best-arm identification (BAI), and provide visualizations in Figure B.2 in Appendix B.
@@
To test hypothesis $R_{\text{THR}}^{0}(a^{\prime})$ at time $t$, we check if $E_{t}(m^{0})$ exceeds the $1/\alpha$ threshold, where entries of $m^{0}$ are given by:
@@
To test hypothesis $R_{\text{THR}}^{1}(a^{\prime})$ at time $t$, we check if $E_{t}(m^{1})$ exceeds the $1/\alpha$ threshold, where entries of $m^{1}$ are defined as:
@@
The vectors $m^{0}$ and $m^{1}$ are the projections of the global minimizer $m_{a}^{*}$ for regions $R_{\text{THR}}^{0}(a^{\prime})$, $R_{\text{THR}}^{0}(a^{\prime})$, respectively, and only involve thresholding the mean of arm $a^{\prime}$ based on $\xi$. Below, we consider a slightly more complicated example, BAI, which has constraints defined between pairs of arms.
@@
For best-arm identification, each arm $a$ has a single hypothesis, as defined in Equation (3). Proposition 3 provides the conditions that define the minima of $E_{t}(m)$ for region $R_{\text{BAI}}(a)$ at time $t$.
@@
Let $R_{\text{BAI}}(a^{\prime})$ be the set defined in Equation (3). Let $\tilde{m}\in[0,1]^{W}$ be the global minimizer of $E_{t}(m)$. Then, $\min_{m\in R_{\text{BAI}}(a^{\prime})}E_{t}(m)$ is obtained at solution $m^{*}$, given by:
$m^{*}_{a}=\begin{cases}\tilde{m}_{a}\quad&\text{if }\tilde{m}_{a}<q\\ q\quad&\text{if }\tilde{m}_{a}\geq q\end{cases},$
@@
$\sum_{a\in W:\tilde{m}_{a}\geq q}\Bigg{(}\gamma(m^{*},a)$
$\sum_{i=1}^{t}\mathbf{1}[A_{t}=a]\frac{2q-X_{i}-\hat{\mu}_{i-1}(a)}{c+(\hat{\mu}_{i-1}(a)-q)(X_{i}-q)}\Bigg{)}=0,$
@@
Figure 2: Log-widths of the confidence sequences across 5000 time steps. Curves represent the average width across 30 simulations.
@@ 5 Empirical Results
[…]
5.1 Synthetic Experiments
Experiments in the Single-Stream Setting
@@ 5.1 Synthetic Experiments
We compare the confidence sequences generated by our test to existing anytime-valid confidence intervals. We use the confidence sequence as a proxy for both the power and growth rate of our test. We compare our approach with four other nonparametric sequential tests that directly apply to bounded means: 1) Chernoff-based sub-Bernoulli (Sub-B) confidence sequences (Howard et al., 2021a, ),888as implemented using the package confseq (Howard et al., 2021b, ). 2) empirical Bernstein (Emp. Bern.) style confidence sequences (Waudby-Smith and Ramdas, 2023, Thm. 2), 3) Hoeffding-based confidence sequences (PrPlH, Waudby-Smith and Ramdas, 2023, Prop. 1), and 4) the hedged capital process confidence sequence (Hedged, Waudby-Smith and Ramdas, 2023, Theorem 3, Eq. 26). For Hedged, we use a grid of 100 points for all simulations. We provide experiment details and additional experiments with nonstandard distributions in Appendix B.
Multi-Streams of Data, Adaptive Sampling   We test multiple streams by 1) introducing adaptive sampling schemes that vary over time, and 2) requiring $\alpha$-level guarantees across multiple parameters of interest. We set $W=4$. In the Bernoulli setting (high variance case), the distribution of arm $a$ is $P_{a}\equiv\text{Bern}(0.15+0.14a)$. In the Beta setting (low variance case), the distribution of arm $a$ is $P_{a}\equiv\text{Beta}(1,\frac{0.85-0.14a}{0.15+0.14a})$. In both cases, $\mu_{a}=0.14a+0.15$. For both THR and BAI, Kano et al.,, 2019 and Jamieson and Nowak,, 2014 (Eq. 2) propose a stopping criterion that uses sequential tests that union bound over each arm. We denote their proposed stopping criterion as Base, and additionally test the two best-performing tests/confidence sequences (Sub-B, Hedged) in the single-arm case by using a similar union-bound approach. We refer to Appendix B for the exact adaptive sampling algorithm used for each bandit identification problem, and further experiment results, including simulations with nonstandard data distributions.
Type Stopping Rule Stopping Time
PEAK 725.81 $\pm$ 218.49
Beta Base 4534.31 $\pm$ 482.94
Sub-B 1618.32 $\pm$ 229.90
Hedged 479.16 $\pm$ 205.97
PEAK 1678.08$\pm$ 666.39
Bern Base 4795.51 $\pm$ 1101.09
Sub-B 1686.81 $\pm$ 456.08
Hedged 2241.18 $\pm$ 1092.49
Table 1: Stopping Times for THR. Stopping times represent the first time that we have classified all 4 arms as above or below $\xi=0.5$. We report averaged stopping times and one standard error over 100 simulated sample paths. Across all simulations, each stopping time results in the correct conclusion (i.e., rejects the incorrect hypothesis).
@@
Discussion of Simulation Results   In the single-arm case, Figure 1 plots the log-width of the confidence sequences across time. The log-width of our confidence interval suggests that PEAK provides a robust testing procedure, even in the single-arm case. Across any combination of variance settings and symmetry, our method performs the second-best across all tested approaches. In the high variance Bernoulli setting, Sub-B performs the best, as the sub-Bernoulli condition is tight with respect to the tested distribution. Compared to all other approaches, our test is closest in performance, corroborating the worst-case optimality results of Lemma 3. For the low-variance case, Hedged outperforms all other methods, but is closely followed by Emp. Bern. and PEAK. We additionally note that PEAK outperforms all benchmarks under mixture distributions as illustrated in Figure B.1.
In the multi-armed case, our test performs comparably to the state-of-the-art testing-by-betting method, Hedged, and outperforms the proposed stopping criterion and Sub-B. For both THR and BAI, the stopping times for PEAK are either the lowest or second lowest on average empirically. For THR (Table B.2), the performance of our approach closely mirrors its performance in the single-arm case, where PEAK performs the second-best for both high and low variance settings, and Hedged and Sub-B perform the best for Beta (low variance) and Bernoulli (high variance) arms, respectively. For BAI, the testing-by-betting methods (Hedged and PEAK) drastically outperform Base and Sub-B as a stopping criterion. For the high variance, Bernoulli setting, PEAK outperforms Hedged, while Hedged outperforms PEAK in the low variance Beta setting.
Runtime   To stress the key benefits of our approach, we test the best-arm hypothesis in the Bernoulli setting across a fixed $T=2000$ time horizon (one test every two time steps), with no stopping, and report the average runtime for the fixed-horizon test over 100 simulations in Table 3. The runtimes for PEAK and Hedged are only comparable when the grid size is 100 (i.e., $10^{-2}$ fidelity) for Hedged. For larger grid sizes, such as 500 evenly spaced points, the runtime of Hedged is at least double that of PEAK, despite having far worse fidelity. The union-bound approach of using confidence intervals with the Hedged betting scheme scales linearly with the fidelity of the grid and becomes exponential if one were to average across arms, as we do with PEAK. This demonstrates that even with a relatively small number of streams, our approach drastically improves upon the computational requirements of existing betting schemes, such as Hedged.